OpenStudy (anonymous):

How would you integrate pi(16sin^2(x)) dx?

OpenStudy (mr.math):

Use \(\sin^2(x)=\large\frac{1- \cos(2x)}{2}\).

OpenStudy (anonymous):

Can I pull out the 16pi?

OpenStudy (mr.math):

Yes.

OpenStudy (anonymous):

Okay I did and I got 4pi^2..is that right?

OpenStudy (anonymous):

oh..end points are 0-pi/2

OpenStudy (anonymous):

0 to pi/2..sorry!

OpenStudy (mr.math):

Correct!

OpenStudy (anonymous):

Thank you very much!! :))

OpenStudy (mr.math):

Glad to help.